OpenStudy (anonymous):

@jim_thompson5910 and @amistre64 please help me :( does any one know how to do this? list the first 5 sequence defined by {-3/2(-1/2)^n-1

OpenStudy (amistre64):

what are the answers for n=1,2,3,4,5

OpenStudy (amistre64):

or at least thats how im interpreting it

OpenStudy (amistre64):

first 5 terms, not sure what first 5 seqs means

OpenStudy (anonymous):

i think its the first answers you get when you put n as any number right

OpenStudy (amistre64):

well, when n equals the first 5 positive integers yes

jimthompson5910 (jim_thompson5910):

I'm guessing it is \[\Large -\frac{3}{2}\left(-\frac{1}{2}\right)^{n-1}\] right?

OpenStudy (anonymous):

so do i plug n as any number to that equation their

OpenStudy (amistre64):

n=1 gives 1st term n=2 gives 2nd term n=3 gives 3rd term ... etc

jimthompson5910 (jim_thompson5910):

no, start off by plugging in n = 1

OpenStudy (amistre64):

2*2^(n-1) = 2^n might simplify the bottom

OpenStudy (amistre64):

and the top is -3*(+-1)

OpenStudy (anonymous):

ok then

OpenStudy (amistre64):

\[\frac{-3}{(-2)^n}\]

OpenStudy (amistre64):

err, might have a wrong sign on that thing

OpenStudy (amistre64):

for a dbl chk .... http://www.wolframalpha.com/input/?i=3%2F%28-2%29%5En

OpenStudy (anonymous):

ok ill check thanks for your help:)